# Scaling Laws for Floating–Point Quantization Training

### v1

@@ 4.1 The Unified Scaling Law Formation
Table 2: Fitted hyper-parameters and their values in our proposed unified scaling law for floating-point quantization training.
Furthermore, we evaluate our scaling law to predict the losses of 1.2B LLMs with different low-precision settings and trained tokens (which are viewed as our validation models that are not used in calculating hyper-parameters of our scaling law). The consistently accurate fitting results demonstrate that our scaling law of floating-point quantization training functions well in larger model sizes.
@@
Figure 14: The fitting results of our scaling law for floating-point quantization training. Data point size is proportional to $D$. The star points (1.2B models) are our validation.
4.2 Implication-1: Optimal Float Layout Analysis
$M_{\text{opt}}=\frac{\nu P}{\delta+\nu}-0.5.$ (26)
More details are in Appendix A. The corresponding loss scaling equation incorporates this optimal layout:
$L(N,D,P,B)=\frac{n}{N^{\alpha}}+\frac{d}{D^{\beta}}+\epsilon+\frac{D^{\beta}}{N^{\alpha}}\cdot\frac{\log_{2}B}{\gamma_{\rho}P^{\delta+\nu}}.$ (27)
where
Figure 15 visualizes the predictive performance for different $P$, demonstrating the effectiveness of the derived layout in preserving performance under varying precisions. The optimal float layouts of FP4, FP8, and FP16 are E2M1, E4M3, and E8M7 (BF16), respectively.
Figure 15: The optimal float layouts of different bit widths.
4.3 Implication-2: Critical Data Size for Optimal Performance
Figure 16: Variation of loss with data size under different floating-point quantization settings.
From Eq. (25) and Fig 16, we can observe that there are two factors containing $D$ that have opposite impacts on LLM loss. Intuitively, it implies that there is an optimal data size under a certain float-pointing quantization setting. The determination of critical data size ($D_{crit}$) stands as a critical juncture within the quantized training regimen. Upon exceeding the threshold of $D_{crit}$ with the training dataset, any additional data introduction negatively impacts the model efficacy, manifesting in a rise in validation loss instead of a decline. A comprehensive derivation for the estimation of $D_{crit}$ is delineated in Appendix B:
4.4 Implication-3: Compute-Optimality with Fixed Configurations
4.4.1 Fixed Data Size $D$
For a fixed $D$, the critical precision $P_{\text{crit}}$ minimizes the loss while accounting for computational constraints. From Appendix C.1, $P_{\text{crit}}$ is expressed as:
$P_{\text{crit}}(D)=\left(\gamma_{D}D^{\beta}\log_{2}B\right)^{\frac{1}{\delta+\nu}},$ (30)
where
$\gamma_{D}$ $=\frac{\delta+\nu-\alpha}{n\alpha\gamma_{\rho}},$ (31)
which consolidates the relationships between model precision and compute efficiency. As illustrated in Figure 17, Eq. (30) suggests that as the amount of training data increases, the most economical precision also correspondingly rises under the constraint of limited computational power. This also suggests that we can adopt such a quantization strategy: in the early stage of training, employ aggressive quantization strategies such as FP8-E4M3, or even FP4-E2M1 which may be available in the future hardware, so as to quickly converge the model to a better level. Subsequently, as the data volume as well as the “knowledge intensity” further increase, gradually enhance the training precision to BF16, or even FP32, in order to maintain the optimal cost-effectiveness of training.
4.4.2 Fixed Model Size $N$
For a fixed $N$, the critical precision depends on balancing the compute resources and maintaining the required training data size. The corresponding scaling law and derivations are detailed in Appendix C.2. We have:
$P_{\text{crit}}(N)$ $=\left[\gamma_{N}\left(\frac{C}{k}\right)^{2\beta}N^{-(\alpha+2\beta)}\log_{2}B\right]^{\frac{1}{\delta+\nu+2\beta}}.$ (32)
$P^{\delta+\nu+2\beta}N^{\alpha+2\beta}=\text{Constant}.$ (33)
4.4.3 Minimization over $N$, $D$, $P$ with Fixed Compute
Through a joint analysis of the impacts of $N$, $D$, and $P$ on the final validation loss, the relationship between cost-effective precision and expected compute budget can be obtained:
$P^{(\delta+\nu)\frac{\alpha+\beta}{\beta}+\alpha}=\lambda\left(\gamma_{D}\log_{2}B\right)^{\frac{\alpha+\beta}{\beta}}\left(\frac{C}{k}\right)^{\alpha},$ (34)
where
$\lambda=\frac{d\beta}{n\alpha}\cdot\frac{\delta+\nu-\alpha}{\delta+\nu+\beta}.$ (35)
In Appendix C.3, we present more detailed derivation processes. Based on the parameters fitted from our experimental data presented in Table 2, with the block size ($B$) set to 128 and $k=6/16$, as illustrated in Figure 18, when the total compute budget is in the range of ($10^{21}$, $10^{31}$) floating-point operations, the optimal cost-performance ratio precision is found to lie between 4 and 8 bits. This finding implies that training larger models with lower precision and utilizing less data yields a more cost-effective approach. Developers could rely on our implications from our scaling law to decide their optimal float-pointing quantization settings.
5 Related Work
5.1 Scaling Law of LLMs
Due to the extremely large cost of resource and time for LLM training, discovering appropriate scaling laws to accurately predict LLM capabilities under different parameters is essential for product-level training. Kaplan et al. (2020) gives the classic form of the scaling law and concludes that the performance penalty depends predictably on the ratio $N^{0.74}/D$. Hoffmann et al. (2022) models the final loss as a parametric function of the count of model parameters and the number of trained tokens, i.e. $E+A/N^{\alpha}+B/D^{\beta}$. Bahri et al. (2024) and Lin et al. (2024) theoretical analysis of how loss scales with the size of the training dataset and the number of parameters in a power-law manner. Previous work (Dettmers & Zettlemoyer, 2023) explores the scaling laws of LLMs with different bit precisions. Recently, Kumar et al. (2024) focuses on the impact of model parameters and data volume with low precision, highlighting the possible negative impacts of more trained tokens in low-precision LLM training and serving. Other recent work (Ouyang et al., 2024) also investigate the correlations of integer quantization and LLM performance. Nevertheless, it is still not that particularly clear to conclude the scaling law for floating-point quantization training with respect to the specific selection of the exponent, mantissa, and block size of scaling factors.
5.2 Quantization of LLMs
The quantization technique of large language models (Lang et al., 2024; Shen et al., 2024) has received widespread attention. Xiao et al. (2023) reduces the accuracy loss during quantization by smoothing the distribution of activations and weights. Dettmers et al. (2024) combines Quantization-Aware Training and LoRA methods to implement an efficient fine-tuning method. Egiazarian et al. (2024) explores techniques for compressing large language models at very low bit rates, and Behdin et al. (2023) proposes a framework that allows each layer to be quantized independently. Although previous work (Zhang et al., 2023; Yoshida, 2023) have studied the impact of exponent, mantissa and blocksize on the quantification of LLMs, the comprehensive impact of these indicators has not been systematically studied and summarized.
6 Conclusion, Limitation, and Future Work
In this work, we propose our scaling law for floating-point quantization training, which functions satisfactorily as a precise guidance of future low-precision LLM training. The key factors of the data size (D), model size (N), exponent (E), mantissa (M), and block size of scaling factors (B) are carefully considered in our scaling law throughout several hundred of experiments with various precision and model settings. Besides the scaling law, we also discover some insightful implications that could instruct and enhance future floating-point quantization training in LLMs. We hope our findings could shed light on better low-prediction LLM training to facilitate the LLM community.
In the future, we will verify our scaling laws for floating-point quantization training under larger model sizes and data sizes. Currently, our explorations are conducted based on the classical Transformer architecture. Whether our scaling laws could also be smoothly applied to LLMs of other architectures (e.g., Mamba series (Dao & Gu, 2024)) is worth confirming. Moreover, our experiments focus on the classical floating-point quantization strategies, while other new-proposed low-bit LLM quantization methods should also be covered in the future.
Appendix A Optimal Float Layout
@@
Appendix B Critical Data Size
For floating-point quantization training, over-training may occur, that is, when the amount of training data exceeds the critical value, the loss will increase instead. Given exponent bits (E), mantissa bits (M), block size (B), and number of model parameters (N), we aim to find the critical data size before over-training. Based on Eq. (23), we derive the expression for $D$ when the partial derivative with respect to $D$ is zero.
@@ Appendix B Critical Data Size
Appendix C Compute-optimality
In order to investigate the optimal precision $P$ under constrained computational budget, we define the computational expenditure associated with floating-point quantization training as follows:
@@ Appendix C Compute-optimality
C.1 Fixed data size $D$
@@
C.2 Fixed model size $N$
@@
C.3 Minimization over $N$, $D$, $P$ with Fixed Compute
Based on the results from Section C.1, C.2 and specifically Eq. (51) with b=0, we proceed to address the system of equations:
@@ C.3 Minimization over $N$, $D$, $P$ with Fixed Compute
Appendix D Ablations

### v3

@@ B.1 Numberical Fits
B.2 Fitting Results of Classical Scaling Laws
We give the fitting results of two classical scaling laws, i.e., the Chinchilla scaling law and the OpenAI scaling law, in Figure 8. We select the Chinchilla scaling law for reference considering its better fitting performance.
(a) Chinchilla.
(b) OpenAI.
Figure 8: Fitting performance of classical scaling laws, with data point size proportional to $D$. Left: Curve based on Chinchilla scaling law shows excellent empirical training loss alignment with predicted losses. Right: Curve based on OpenAI scaling law also demonstrates a good match, though less precise than Chinchilla.
B.3 Fitting Results of the Scaling Laws for Exponent and Mantissa
Figure 9 shows the fitting performance of our scaling law related to the exponent or mantissa, and Figure 10 shows the correlations between essential parameters in Eq. (9).
(a) Exponent.
(b) Mantissa.
Figure 9: The fitting outcomes of our scaling law related to the exponent (left)/mantissa (right). Data point size is proportional to $D$.
Figure 10: The correlations between $\gamma$,$\iota$ in Eq. (9) and $N$,$D$. $\gamma$,$\iota$ can be viewed as functions of $N$,$D$. Data point size is proportional to $D$.
B.4 Fitting Results of the Scaling Laws for Scaling Factor
These figures show different correlations and fitting results related to the scaling factor B in our scaling law.
(a) Channel-wise.
(b) Tensor-wise.
@@ B.4 Fitting Results of the Scaling Laws for Scaling Factor
Figure 13: Fitting results of the scaling law under different sharing strategies of the scaling factor. The size of the data point is proportional to $D$.
Appendix C Optimal Float Layout
@@
Appendix D Critical Data Size
For FP quantization training, overtraining may occur, that is, when the amount of training data exceeds the critical value, the loss will increase instead. Given exponent bits (E), mantissa bits (M), block size (B), and number of model parameters (N), we aim to find the critical data size before over-training. Based on Eq. (16), we derive the expression for $D$ when the partial derivative with respect to $D$ is zero.
@@ Appendix D Critical Data Size
Appendix E Compute-optimality
In order to investigate the optimal precision $P$ under constrained computational budget, we define the computational expenditure associated with FP quantization training as follows:
@@ Appendix E Compute-optimality
E.1 Fixed data size $D$
@@
As illustrated in Figure 14, Eq. (47) suggests that as the amount of training data increases, the most economical precision also correspondingly rises under the constraint of limited computational power.
E.2 Fixed model size $N$
@@
E.3 Minimization over $N$, $D$, $P$ with Fixed Compute
Based on the results from Section E.1, E.2 and specifically Eq. (42) with b=0, we proceed to address the system of equations:
@@ E.3 Minimization over $N$, $D$, $P$ with Fixed Compute
Appendix F Detailed Settings of Experiments
We show the detailed configurations of our experiments as follows: